7) A company pays tax at a rate of 15% on its first $50,000 of income. If it has $60,000 of 7) taxable income and an average tax rate of 18%, what is the marginal tax rate on its last $10,000 of income? A) 25% ?) 18% C) 33% D) 15% 8) According to the statement of cash flows, cash flows from financing could be positive 8) ?f: A) the firm added more debt than it repaid. B) interest rates were low on outstanding debt. C) the firm sold portions of its plant and equipment. D) the firm repaid more debt than it added. 9) Current period depreciation expense is listed 9) A) on the balance sheet. B) on the income statement. C) in the investment section of the cash flow statement. D) on neither the balance sheet nor the income statement; it is a noncash expense. 10) Which of the following is not a problem with EVA? 10) A) EVA cannot be used to measure the profitability of a private company B) EVA assumes that you know the cost of capital C) EVA cannot be used to compare the effectiveness of managers with different amounts of assets under their control D) EVA assumes that the book values of assets are equal to their current worth 11) Which of the following actions could improve a firms current ratio if it is now less than 11) 1.0? A) Paying accounts payable with cash B) Buying inventory on credit C) Converting marketable securities to cash D) Selling inventory at cost 12) Lease obligations are included in certain leverage ratios because leases: 12) A) must be financed through a bank. B) are perpetual obligations. C) represent long-term fixed obligations. D) require the payment of interest. 13) 13) What is the minimum nominal rate of return that you should accept if you require a 496 real rate of return and the rate of inflation is expected to average 3.5% during the investment period? A) 7.50% B) 7.64% C) 8.01% D) 7.36%
Answer #1

7. Total tax on $60,000 X 18% = $ 10,800

Tax on First slab $ 50,000 X 15% = $ 7,500

Differntial Tax = $ 3,300 on additional tax of $ 10,000 is 33%

Correct answer is C

8. If firm adds more debt (raise) then it will increase the cash inflow from financing activities. Hence correct answer is A

9. in the income statement as it is charge to P&L for the year. B

10. Correct answer is C

11. If firm buys inventory on credit it will improve the Current ration from present level

12. C as it is long term fixed obligation which have to be paid and hence should be cover throug the either EBITDA or accrual

13. (1+R)(1+i) = (1+N)

R is real rate, i is inflation and N is nominal rate

(1+4%)(1+3.5) = 1+N

Solving this gives, N = 7.64% hence answer is B